# Enhancing operational wind downscaling capabilities over Canada: Application of a Conditional Wasserstein GAN methodology

### v1

@@ Abstract
Wind downscaling is essential for improving the spatial resolution of weather forecasts, particularly in operational Numerical Weather Prediction (NWP). This study advances wind downscaling by extending the DownGAN framework introduced by Annau et al. to operational datasets from the Global Deterministic Prediction System (GDPS) and High-Resolution Deterministic Prediction System (HRDPS), covering the entire Canadian domain. We enhance the model by incorporating high-resolution static covariates, such as HRDPS-derived topography, into a Conditional Wasserstein Generative Adversarial Network with Gradient Penalty, implemented using a UNET-based generator. Following the DownGAN framework, our methodology integrates low-resolution GDPS forecasts (15 km, 10-day horizon) and high-resolution HRDPS forecasts (2.5 km, 48-hour horizon) with Frequency Separation techniques adapted from computer vision. Through robust training and inference over the Canadian region, we demonstrate the operational scalability of our approach, achieving significant improvements in wind downscaling accuracy. Statistical validation highlights reductions in root mean square error (RMSE) and log spectral distance (LSD) metrics compared to the original DownGAN. High-resolution conditioning covariates and Frequency Separation strategies prove instrumental in enhancing model performance. This work underscores the potential for extending high-resolution wind forecasts beyond the 48-hour horizon, bridging the gap to the 10-day low resolution global forecast window.
@@ 1 Introduction
Accurate high-resolution wind forecasts are crucial for several application sectors such as aviation [1], wind energy [2], power infrastructure [3], public safety [4], wildfire management [5], avalanche warnings [6], winter sports [7], air quality [8] and water quality management [9]. Typically, wind forecasts are obtained from numerical weather predictions (NWPs), starting with a global forecast model followed by a limited-area forecast model which takes the lower-resolution model as boundary conditions in a procedure called dynamical downscaling. In Canada, high-resolution numerical prediction models have a limited forecast horizon of 48 hours due to computational constraints. An applied perspective of the work presented in this paper is to extend high-resolution wind forecast services from 48 hours to the same 10 days forecast horizon as the Canadian global forecast model.
@@
More recently, several contributions have demonstrated the potential of Artificial Intelligence (AI) based downscaling [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], applying successfully techniques originally developed for image processing and computer vision to weather data. In the field of image processing, the image super-resolution (or zooming) problem consists of taking a coarser image as input and producing a finer image at higher resolution. [^1] Confusingly, what is called upsampling (downsampling) in the field of image processing is called downscaling (upscaling) in meteorology.
@@
Earlier methods were achieving an enhancement factor of 2 to 4 times, but with the advent of deep learning, an enhancement factor of 8 and more is now possible. Generative AI in particular has proven to be a powerful tool not only to artificially produce realistic-looking images and videos from a random seed, but also to condition the generative AI on coarser or partially missing data (called inpainting).
@@
AI-based techniques for downscaling weather and climate fields can be classified in an increasing order of complexity and recency. At the end of spectrum, ensemble decision trees techniques such random forests and extreme gradient boosting are the most similar to MLR and the easiest to implement. Other classical machine learning techniques such as support vector machine and shallow artificial neural networks with a single hidden layer can also fit in that category. Methods based on Convolutional Neural Networks (CNNs) have proved to be both popular and powerful. Generative Adversarial Networks (GANs) are the first type of generative AI allowing to sample an ensemble of possible realizations [15]. More recently, latent diffusion techniques have emerged as an interesting alternative to GAN. Finally, methods based on vision transformers and in particular those that consist of fine-tuning a weather foundation model are emerging.
@@
In the literature, these methods have been applied to a variety of weather fields such as temperature, wind speed or wind vectors, and precipitation. Many of the published work focus on one or several small areas as a proof-of-concept, and few are used in an operational setting. It is important to note that similar AI techniques can be used either for downscaling weather forecasts, climate projections or reanalysis data, with the main difference in the choice and resolution of source and target data. Moreover, with a general downscaling model , there is no conceptual difference between downscaling to a finer grid or downscaling to a point location that could be either an existing weather station or any location of interest.
@@
The key contributions of this work are as follows: 1) We extend the DownGAN framework introduced by [15] by applying its methods to operational Numerical Weather Prediction (NWP) datasets, specifically GDPS and HRDPS across the entire Canadian domain, expanding beyond the regional focus of the original study. 2) We enhance the DownGAN model, a Wasserstein GAN with Gradient Penalty (WGAN-GP) framework for wind downscaling, by introducing high-resolution static covariates, such as topography from the HRDPS domain, as conditioning inputs within the generator architecture, implemented using a UNET. We refer this modified framework as a Conditional WGAN-GP. 3) We propose a robust training and inference schema that allows the model to effectively learn wind patterns and generate predictions over the entire Canadian region. This development ensures the scalability and operational relevance of the methodology. 4) We perform statistical validation for large-scale wind downscaling across the Canadian domain, identifying limitations such as the ineffectiveness of Annau’s partial frequency separation technique in this context. However, Frequency Separation methods adapted from computer vision and used by DownGAN show strong performance. The validation highlights the importance of high-resolution conditioning covariates and the utility of Frequency Separation in this context.
@@ 2 Dataset
2.1 Global Deterministic Prediction System
The Global Deterministic Prediction System (GDPS) [26] is ECCC’s operational long-range forecast model. It represents the global configuration of the Global Environmental Multiscale (GEM) model [27] and it is run twice a day (at 00 UTC and 12 UTC), producing 10-day forecasts at a 15-km nominal resolution.
2.2 High-Resolution Deterministic Prediction System
@@ 2.3 Predictors and Predictands
High-resolution zonal (East-West) and meridional (North-South) surface (10 m) wind components from HRDPS are used as the target variables. By predicting both components of the wind vector, we can derive both wind speed and direction at the resolution of HRDPS. Low-resolution zonal and meridional surface wind from GDPS are taken as predictors along with surface temperature, as well as wind, temperature and vertical motion at 546 hPa. In contrast with [15], high-resolution geophysical fields are added as predictors. These static high-resolution covariates can be used as predictors since these variables are constant at the time-scale of weather prediction. The topography field is particularly important in mountainous regions where wind channeling is predominant. The land-water mask not only provides an indication of changing surface roughness between the smoother water and other land, but can also be a driver of local winds such as sea breezes and lake breezes. Finally, the surface roughness variable modulates the speed of the wind and changing surface roughness can drive turbulence in the planet boundary layer. Table LABEL:table:vars_description describes the meteorological variables used for training the AI downscaling model.
Table 1: Description of Variables for GDPS and HRDPS
@@
True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s].
True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s].
Air temperature at the surface (1.5 m) [°C].
Air temperature vertically interpolated at 546 hPa [°C].
True geographical West-East (zonal) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
True geographical South-North (meridional) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
Vertical motion vertically interpolated at 546 hPa [Pa/s].
@@
True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s]. Denoted by u10.
True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s]. Denoted by v10.
@@
Model orography [m].
Water/land mask [fraction].
Roughness length [m].
@@ 2.4 Data Pre-processing
The map projection of the GDPS outputs is a Yin-Yang grid with a different rotated latitude-longitude map projection than the one of the HRDPS grid. We regrid the GDPS grid to the HRDPS grid using nearest neighbor interpolation. The interpolated data is then reduced by a factor of 8 (20-km nominal resolution) as the input of the AI downscaling method. Fig. 1 illustrates the respective domains of GDPS (cropped over North-America) and of HRDPS.
@@ 3 Methods
In this study, we adopt the deep learning framework for downscaling wind components proposed by [15]. This framework is based on Wasserstein Generative Adversarial Networks with Gradient Penalty (WGAN-GP) [29] and incorporates elements from Generative Adversarial Networks for image super-resolution (SRGAN) [30] as a baseline for the downscaling task. In the following sections, we elaborate on the application of WGANs for the downscaling problem, describe the model architecture, and introduce our approach to integrating high-resolution static covariates into the training process using UNETs [31].
@@ 3.1 Conditional WGAN-GP for Downscaling
Given a training set consisting of dynamic low-resolution (GDPS) and high-resolution (HRDPS) forecast hours, denoted as ${\bf x}$ and ${\bf y}$, respectively. A conditional WGAN-GP, conditioned on the covariates ${\bf y_{cov}}$ in the HRDPS domain, aims to minimize the following losses for the critic $\mathcal{L}_{C}$ and generator $\mathcal{L}_{G}$ [15, 29, 32]:
@@
$\mathcal{L}_{G}=-\gamma\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]+\alpha\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}}))$ (2)
In this formulation, $C$ denotes a real-valued function known as the Critic, which is defined on the HRDPS domain. The function $G$ represents the Generator, which maps low-resolution GDPS to the HRDPS domain. The symbol $\mathbb{P}_{g}$ denotes the conditional distribution of the generated downscaled GDPS data given the covariates ${\bf y_{cov}}$, while $\mathbb{P}_{r}$ represents the distribution of the HRDPS points. The expression $\mathbb{E}_{{\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}}[(\|\nabla_{\hat{y}}C(\hat{y})\|_{2}-1)^{2}]$ represents the gradient penalty, which is incorporated into the formulation to ensure valid Wasserstein distances by enforcing 1-Lipschitz continuity for the Critic. Here, ${\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}$ denotes a linearly interpolated generated value defined by ${\bf\hat{y}}=\epsilon{\bf y}+(1-\epsilon)G({\bf x}\mid{\bf y_{cov}})$, where $\epsilon\sim U[0,1]$. The scalars $\lambda$, $\gamma$, and $\alpha$ are regularization hyperparameters: $\lambda$ controls the gradient penalty term, $\gamma$ the adversarial loss term, $\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$ and $\alpha$ the content loss term, $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}}))$. All expectations in the equations above are computed across all dimensions of ${\bf x}$ and ${\bf y}$. Specifically, ${\bf x}$ is a tensor of size $B\times C_{l}\times W_{l}\times H_{l}$, and ${\bf y}$ is a tensor of size $B\times C_{h}\times W_{h}\times H_{h}$. Here, $B$ represents the batch size, $W$ and $H$ denote the spatial width and height, respectively, while $C$ indicates the number of input channels (variates). The subscripts $l$ and $h$ refer to the GDPS and HRDPS datasets, respectively. Additionally, ${\bf y_{cov}}$ is a tensor with dimensions $C_{cov}\times W_{l}\times H_{l}$, where values are replicated across the batch size $B$. This replication is due to the covariates being static and independent of time. It is important to note that if ${\bf y_{cov}}$ is not provided, the WGAN-GP model reduces to the formulation presented in [15].
@@
3.1.1 Frequency Separation
The core idea of the frequency separation approach is to decompose the low- and high-frequency components and treat them differently during WGAN-GP training [15, 34], with the aim of enhancing the quality of downscaled wind components. This allows the adversarial training to focus specifically on generating the high-frequency components, which are crucial for capturing fine details. In this approach, the high-frequency components are derived as ${\bf y}_{h}={\bf y}-H({\bf y})$ for the HRDPS data, and $G({\bf x}\mid{\bf y_{cov}})_{h}=G({\bf x}\mid{\bf y_{cov}})-H(G({\bf x}\mid{\bf y_{cov}}))$ for the generated downscaled data, where $H$ represents a low-pass filter. In this case, during WGAN-GP training, the Critic will focus on optimizing the loss based only in the high frequency components ${\bf y}_{h}$ and $(G({\bf x}\mid{\bf y_{cov}})_{h}$ instead of ${\bf y}$ and $(G({\bf x}\mid{\bf y_{cov}}))$. On the other side the Generator delegates the generation of high frequency components to the adversarial component $-\mathbb{E}_{G({\bf x})_{h}\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))_{h}]$ and the low frequency components to the content loss $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}(H({\bf y}),H(G({\bf x}\mid{\bf y_{cov}})))$. In [15] is proposed the partial frequency separation, this approach is similar than before but the adversarial loss is used in its original formulation, i.e., $-\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$

### v3

@@ Abstract
Downscaling based on Artificial Intelligence (AI) has been demonstrated to be a powerful tool for improving the spatial resolution of weather forecasts. This study advances wind downscaling by extending the DownGAN framework introduced by Annau et al. to operational Canadian Numerical Weather Predictions (NWP), namely from the Global Deterministic Prediction System (GDPS) to the domain of the High-Resolution Deterministic Prediction System (HRDPS). We enhance the AI model by incorporating high-resolution static covariates, such as topography on the HRDPS grid, into a Conditional Wasserstein Generative Adversarial Network with Gradient Penalty, implemented using a UNET-based generator. Following the DownGAN framework, our methodology integrates low-resolution GDPS forecasts (15 km, 10-day horizon) and high-resolution HRDPS forecasts (2.5 km, 48-hour horizon) with Frequency Separation techniques adapted from computer vision. Through robust training and inference over the whole Canadian region, we demonstrate the operational scalability of our approach, while achieving significant improvements in wind downscaling accuracy. Statistical validation highlights reduction in root mean square error (RMSE) and log spectral distance (LSD) metrics compared to the original DownGAN. High-resolution conditioning covariates and Frequency Separation strategies prove instrumental in enhancing model performance. This work underscores the potential for extending high-resolution wind forecasts beyond the 48-hour horizon, bridging the gap to the 10-day low resolution global forecast window.
@@ 1 Introduction
Accurate high-resolution wind forecasts are crucial for several application sectors such as aviation [1], wind energy [2], power infrastructure [3], public safety [4], wildfire management [5], avalanche warnings [6], winter sports [7], air quality [8] and water quality management [9]. Typically, high-resolution wind forecasts are obtained from numerical weather predictions (NWPs), starting with a global forecast model followed by a limited-area forecast model which takes the lower-resolution model as boundary conditions in a procedure called dynamical downscaling. In Canada, high-resolution numerical prediction models have a limited forecast horizon of 48 hours due to computational constraints. An applied perspective of the work presented in this paper is to extend high-resolution wind forecast services from 48 hours to the same 10 days forecast horizon as the Canadian global forecast model.
@@
More recently, several contributions have demonstrated the potential of Artificial Intelligence (AI) based downscaling [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], applying successfully techniques originally developed for image processing and computer vision to weather data.
@@
In the field of image processing, the image super-resolution (or zooming) problem consists of taking a coarser image as input and producing a finer image at higher resolution. [^1] Confusingly, what is called upsampling (downsampling) in the field of image processing is called downscaling (upscaling) in meteorology. Earlier image super-resolution methods were achieving an enhancement factor of 2 to 4 times, but with the advent of deep learning, an enhancement factor of 8 and more is now possible. Generative AI in particular has proven to be a powerful tool not only to artificially produce realistic-looking images and videos from a random seed, but also to condition the generative AI on coarser or partially missing data (called inpainting).
@@
AI-based techniques for downscaling weather and climate fields can be classified in an increasing order of complexity and recency. At the end of spectrum, ensemble decision trees [17] techniques such random forests and extreme gradient boosting are the most similar to MLR and the easiest to implement. Other classical machine learning techniques such as support vector machine and shallow artificial neural networks with a single hidden layer can also fit in that category. Methods based on Convolutional Neural Networks (CNNs) [18, 25, 26], have proved to be both popular and powerful. Generative Adversarial Networks (GANs) are the first type of generative AI allowing to sample an ensemble of possible realizations [15, 24, 27]. More recently, latent diffusion techniques have emerged as an interesting alternative to GAN [28]. Finally, methods based on vision transformers and in particular those that consist of fine-tuning a weather foundation model are emerging.
@@
In the literature, these methods have been applied to a variety of weather fields such as temperature, wind speed or wind vectors, and precipitation. Many of the published works focus on one or several small areas as a proof-of-concept, and few are used in an operational setting. It is important to note that similar AI techniques can be used either for downscaling weather forecasts, climate projections or reanalysis data, with the main difference in the choice and resolution of source and target data. Moreover, with a general downscaling model, there is no conceptual difference between downscaling to a finer grid or downscaling to a point location that could be either an existing weather station or any location of interest.
@@
The key contributions of this work are as follows: 1) We extend the DownGAN framework introduced by [15] by applying its methods to operational Numerical Weather Prediction (NWP) datasets, specifically GDPS and HRDPS across the entire Canadian domain, expanding beyond the regional focus of the original study. 2) We enhance the DownGAN model, a Wasserstein GAN with Gradient Penalty (WGAN-GP) framework for wind downscaling, by introducing high-resolution static covariates, such as topography from the HRDPS domain, as conditioning inputs within the generator architecture, implemented using a UNET. We refer this modified framework as a Conditional WGAN-GP. 3) We propose a robust training and inference schema that allows the model to effectively learn wind patterns and generate predictions over the entire Canadian region. This development ensures the scalability and operational relevance of the methodology. 4) We perform statistical validation for large-scale wind downscaling across the Canadian domain, identifying limitations such as the ineffectiveness of Annau’s partial frequency separation technique in this context. However, Frequency Separation methods adapted from computer vision and used by DownGAN show strong performance. The validation highlights the importance of high-resolution conditioning covariates and the utility of Frequency Separation for downscaling.
@@ 2 Dataset
2.1 Numerical Weather Predictions
Recall that the goal of this study is to extend the forecast horizon of the High-Resolution Deterministic Prediction System (HRDPS) by downscaling the Global Deterministic Prediction System (GDPS). The Global Deterministic Prediction System (GDPS) [29] is ECCC’s operational long-range forecast model. It represents the global configuration of the Global Environmental Multiscale (GEM) model [30] and it is run twice a day (at 00 UTC and 12 UTC), producing 10-day forecasts at a 15-km nominal resolution.
@@ 2.2 Predictors and Predictands
High-resolution zonal (East-West) and meridional (North-South) surface (10 m) wind components from HRDPS are used as the target variables. By predicting both components of the wind vector, we can derive both wind speed and direction at the resolution of HRDPS. Low-resolution zonal and meridional surface wind from GDPS are taken as predictors along with surface temperature, as well as wind, temperature and vertical motion at 546 hPa. While we desired to stay close to the choices of predictors of [15], we did not include the Convective Available Energy Potential (CAPE) as it was not readily available in GDPS with an hourly resolution. We instead chose surface and mid-troposphere variables as they are expected to be important drivers of surface winds. However, we did not perform any variable selection experiment. In contrast with [15], the following three high-resolution geophysical fields are added as predictors: model orography, land-water mask and surface roughness length. These static high-resolution covariates can be used as predictors since these variables are constant at the time-scale of weather prediction. The model orography field is particularly important in mountainous regions where wind channeling is predominant. The land-water mask not only provides an indication of changing surface roughness between the smoother water and other land, but can also be a driver of local winds such as sea breezes and lake breezes. Finally, the surface roughness variable modulates the speed of the wind, and changes in this covariate can drive turbulence in the planetary boundary layer. Table LABEL:table:vars_description describes the meteorological variables used for training the AI downscaling model.
Table 1: Description of predictors and predictands for the downscaling experiments. The HRDPS static covariates are extra high-resolution predictors used in the Conditional WGAN-GP, whereas the DownGAN baseline we implemented only uses GDPS predictors.
@@
$\mbox{U}_{\mbox{surf}}$ True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s].
$\mbox{V}_{\mbox{surf}}$ True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s].
$\mbox{T}_{\mbox{surf}}$ Air temperature at the surface (1.5 m) [°C].
$\mbox{T}_{546}$ Air temperature vertically interpolated at 546 hPa [°C].
$\mbox{U}_{546}$ True geographical West-East (zonal) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
$\mbox{V}_{546}$ True geographical South-North (meridional) component of the horizontal wind vertically interpolated at 546 hPa [m/s].
$\mbox{W}_{546}$ Vertical motion vertically interpolated at 546 hPa [Pa/s].
@@
u10 True geographical West-East (zonal) component of the horizontal wind at the surface (10 m) [m/s].
v10 True geographical South-North (meridional) component of the horizontal wind at the surface (10 m) [m/s].
@@
me Model orography [m].
mg Water/land mask [fraction].
z0 Roughness length [m].
@@ 2.3 Data pre-processing
The map projection of the GDPS outputs is a Yin-Yang grid with a different rotated latitude-longitude map projection than the one of the HRDPS grid. To align the grids, we project [^2] The regridding process was done using the xESMF Python Library [32]. the GDPS grid to the HRDPS grid using nearest neighbor interpolation. The interpolated data is then reduced by a factor of 8 (20-km nominal resolution) as the input of the AI downscaling method. Fig. 1 illustrates the respective domains of GDPS (cropped over North-America) and of HRDPS.
Figure 1: Example of zonal and meridional winds on the GDPS and HRDPS grids. Note that the GDPS grid was cropped over North America. The domain of HRDPS grid is indicated by a black rectangle.
@@ 3 Methods
In this study, we adapt the deep learning framework for downscaling wind components proposed by [15]. This framework is based on Wasserstein Generative Adversarial Networks with Gradient Penalty (WGAN-GP) [33] and incorporates elements from Generative Adversarial Networks for image super-resolution (SRGAN) [34] as a baseline for the downscaling task. In the following sections, we elaborate on the application of WGANs for the downscaling problem, describe the model architecture, and introduce our approach for integrating high-resolution static covariates into the training process using UNETs [35].
@@ 3.1 Conditional WGAN-GP for Downscaling
We consider a training set consisting of dynamic low-resolution predictors, such as GDPS fields, and high-resolution predictands, such as HRDPS fields, with matched forecast hours, which we denote as ${\bf x}$ and ${\bf y}$, respectively. A conditional WGAN-GP, conditioned on the covariates ${\bf y_{cov}}$ in the HRDPS domain, aims to minimize the following losses for, respectively, the critic $\mathcal{L}_{C}$ and generator $\mathcal{L}_{G}$ [15, 33, 36]:
@@
and
$\mathcal{L}_{G}=-\gamma\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]+\alpha\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}})).$ (2)
In this formulation, $C$ denotes a real-valued function known as the Critic, which is defined on the high-resolution domain. The function $G$ represents the Generator, which maps low-resolution predictors to the high-resolution predictands. The symbol $\mathbb{P}_{g}$ denotes the conditional distribution of the generated downscaled data given the high-resolution static covariates ${\bf y_{cov}}$, while $\mathbb{P}_{r}$ represents the distribution of the high-resolution predictands. The expression $\mathbb{E}_{{\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}}[(\|\nabla_{\hat{y}}C(\hat{y})\|_{2}-1)^{2}]$ represents the gradient penalty, which is incorporated into the formulation to ensure valid Wasserstein distances by enforcing 1-Lipschitz continuity for the Critic. Here, ${\bf\hat{y}}\sim\mathbb{P}_{\hat{y}}$ denotes a linearly interpolated generated value defined by ${\bf\hat{y}}=\epsilon{\bf y}+(1-\epsilon)G({\bf x}\mid{\bf y_{cov}})$, where $\epsilon\sim U[0,1]$. The scalars $\lambda$, $\gamma$, and $\alpha$ are regularization hyperparameters: $\lambda$ controls the gradient penalty term, $\gamma$ the adversarial loss term, $\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$ and $\alpha$ the content loss term, $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}({\bf y},G({\bf x}\mid{\bf y_{cov}}))$. All expectations in the equations above are computed across all dimensions of ${\bf x}$ and ${\bf y}$. Specifically, ${\bf x}$ is a tensor of size $B\times C_{l}\times W_{l}\times H_{l}$, and ${\bf y}$ is a tensor of size $B\times C_{h}\times W_{h}\times H_{h}$. Here, $B$ represents the batch size, $W$ and $H$ denote the spatial width and height, respectively, while $C$ indicates the number of input channels (variates). The subscripts $l$ and $h$ refer to the high- and low-resolution datasets, respectively. Additionally, ${\bf y_{cov}}$ is a tensor with dimensions $C_{cov}\times W_{l}\times H_{l}$, where values are replicated across the batch size $B$. This replication is due to the covariates being static and independent of time. It is important to note that if ${\bf y_{cov}}$ is not provided, the WGAN-GP model reduces to the formulation presented in [15].
@@
3.2 Frequency Separation
The core idea of the frequency separation approach is to decompose the low- and high-frequency components and treat them differently during WGAN-GP training [15, 38], with the aim of enhancing the quality of downscaled wind components. This allows the adversarial training to focus specifically on generating the high-frequency components, which are crucial for capturing fine details. In this approach, the high-frequency components are derived as ${\bf y}_{high}={\bf y}-\mathscr{L}({\bf y})$ for the HRDPS data, and $G({\bf x}\mid{\bf y_{cov}})_{high}=G({\bf x}\mid{\bf y_{cov}})-\mathscr{L}(G({\bf x}\mid{\bf y_{cov}}))$ for the generated downscaled data, where $\mathscr{L}$ represents a low-pass filter. In this case, during WGAN-GP training, the Critic will focus on optimizing the loss based only in the high frequency components ${\bf y}_{high}$ and $G({\bf x}\mid{\bf y_{cov}})_{high}$ instead of ${\bf y}$ and $G({\bf x}\mid{\bf y_{cov}})$. On the other side the Generator delegates the generation of high frequency components to the adversarial loss component $-\mathbb{E}_{G({\bf x})_{high}\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}})_{high})]$ and the low frequency components to the content loss $\mathbb{E}_{{\bf y}\sim\mathbb{P}_{r}}l_{c}(\mathscr{L}({\bf y}),\mathscr{L}(G({\bf x}\mid{\bf y_{cov}})))$. [^3] Annau et.al., proposed partial frequency separation [15], this approach is similar to frequency separation, but the adversarial loss is used in its original formulation, i.e., $-\mathbb{E}_{G({\bf x})\sim\mathbb{P}_{g}}[C(G({\bf x}\mid{\bf y_{cov}}))]$.